Menlo Company distributes a single product. The company’s sales and expenses for last month follow: Total Per Unit Sales $ 310,000 $ 20 Variable expenses 217,000 14 Contribution margin 93,000 $ 6 Fixed expenses 76,800 Net operating income $ 16,200 Required: 1. What is the monthly break-even point in unit sales and in dollar sales? 2. Without resorting to computations, what is the total contribution margin at the break-even point? 3-a. How many units would have to be sold each month to attain a target profit of $37,200? 3-b. Verify your answer by preparing a contribution format income statement at the target sales level. 4. Refer to the original data. Compute the company's margin of safety in both dollar and percentage terms. 5. What is the company’s CM ratio? If the company can sell more units thereby increasing sales by $66,000 per month and there is no change in fixed expenses, by how much would you expect monthly net operating income to increase?
Answer #1

1. Break even point(units) =fixed cost/contribution per unit =76800/(20-14) =76800/6 =12800 Break even point(dollars) =fixed